OpenStudy (anonymous):

charile runs at a speed of 3 yards per seconde. About how many miles per hour does charile run?

OpenStudy (anonymous):

plz help any one plz

OpenStudy (anonymous):

hey r u going to help me

OpenStudy (jdoe0001):

hmm how many feet are there in 1 mile?

OpenStudy (anonymous):

okok it is 1760

OpenStudy (anonymous):

hurry plz

OpenStudy (jdoe0001):

ok, so yes, there are 5280 feet per mile 3 yards per feet, thus 5280/3 = 1760 so we know there are 1760 yards per mile we also know that there are 60 minutes in 1 hr \(\bf \cfrac{60min}{1hr}\) and there are 60secs in 1 minute \(\bf \cfrac{60sec}{min}\) so charlie runs 3yards in 1 sec, that is \(\bf \cfrac{3yd}{1sec}\) so we can say that \(\bf \cfrac{3\cancel{yd}}{1\cancel{sec}} \times \cfrac{60\cancel{min}}{1hr} \times \cfrac{60\cancel{sec}}{\cancel{min}} \times \cfrac{1mile}{1760\cancel{yd}}\)

OpenStudy (anonymous):

so thats the answer right

OpenStudy (jdoe0001):

that's how you'd convert it, by using the relative units

OpenStudy (anonymous):

ok wait about one minute tell i write it down ok

OpenStudy (anonymous):

ok wats next

OpenStudy (jdoe0001):

that's all you need to get it :)

OpenStudy (anonymous):

ok well wats the answer than

OpenStudy (jdoe0001):

you get it from that unit conversion

OpenStudy (anonymous):

well wat do i do there thats wat im meaning

OpenStudy (jdoe0001):

you multiply the fractions, \(\bf \cfrac{3yd}{1sec} \times \cfrac{60min}{1hr} \times \cfrac{60sec}{min} \times \cfrac{1mile}{1760yd}\) what do you get?

OpenStudy (anonymous):

ok wat 1 minute

OpenStudy (anonymous):

so its 10800 over 1760 right?

OpenStudy (anonymous):

is it

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

r u sure

OpenStudy (anonymous):

now tell me the truth plz

OpenStudy (anonymous):

come on plz hurry

OpenStudy (anonymous):

type faster yo

OpenStudy (jdoe0001):

heheh, well, it's so, yes \(\bf \bf \cfrac{3\cancel{yd}}{1\cancel{sec}} \times \cfrac{60\cancel{min}}{1hr} \times \cfrac{60\cancel{sec}}{\cancel{min}} \times \cfrac{1mile}{1760\cancel{yd}} = \cfrac{10800mi}{1760hr}\)

OpenStudy (anonymous):

oh ok well thats a no wonder im a straight a studenet and do we simpify that yes or no

OpenStudy (jdoe0001):

well, yes, you'd need to

OpenStudy (anonymous):

ok wat grade r u in

OpenStudy (jdoe0001):

well, keep in mind that we "help", at least I do, we do not give flat answers, so the meaning of "help" is to help you understand how to do the process so the process is to use the relevant units and cancel out accordingly

OpenStudy (anonymous):

ok thx a lot and am i ur fan

OpenStudy (anonymous):

oh yes i am lol:)

OpenStudy (anonymous):

thx bye gtg

hero (hero):

You still have to reduce it

hero (hero):

@magie22, you still have to reduce it.